Question 5 15 pts Consider the following matrices: 2 -21 3 0 9 A= -1 1 1 C=0 -6 0 T 3 5 12 2 5 Find matrix B so that: AB=C
Answer #1

The given problem solve Using matrix inversion.3 og Here, A= 2 -2 -2 l -1 11 -135 C. 0-60 12 25 How, det a 2 -21 -1 35 2215-3) + (-10-3) -14-2-1) 2 = 4 -1373 --6 T 2 4 -2 aGiven that ADC Ba Al 2 12 -3 ܦ 30 6 6- ܘ ][? [ 6 ܟܐ ܐܙ ܘ - 2- 12 2 5 3 35- ܘ3 *. 2) 9 29 -6 -#2 25 8le ܐܐܐ ܟܐ 4 ]2 -/2 - 5 #3